\providecommand{\bysame}{\leavevmode\hbox to3em{\hrulefill}\thinspace}
\providecommand{\MR}{\relax\ifhmode\unskip\space\fi MR }
% \MRhref is called by the amsart/book/proc definition of \MR.
\providecommand{\MRhref}[2]{%
  \href{http://www.ams.org/mathscinet-getitem?mr=#1}{#2}
}
\providecommand{\href}[2]{#2}
\begin{thebibliography}{1}

\bibitem{Apelblat2}
Alexander Apelblat, \emph{Tables of integrals and series}, 1 ed., Harri
  Deutsch, Thun, 1996.

\bibitem{BorweinEM10}
Jonathan~Michael Borwein, David~J. Broadhurst, and Joel Kamnitzer,
  \emph{Central binomial sums, multiple {C}lausen values and {Z}eta values},
  Exp.\ Math. \textbf{10} (2001), no.~1, 25--34. \MR{1821569 (2002k:11105)}

\bibitem{BorweinAM70}
Jonathan~Michael Borwein and Roland Girgensohn, \emph{Evaluations of binomial
  series}, Aequat. Math. \textbf{70} (2005), no.~1--2, 25--36. \MR{2167981}

\bibitem{DavydychevNPB699}
A.~I. Davydychev and M.~Yu. Kalmykov, \emph{Massive {Feynman} diagrams and
  inverse binomial sums}, Nucl.\ Phys.\ B \textbf{699} (2004), no.~1--2, 3--64.
  \MR{2095786}

\bibitem{LehmerAMM92}
Derrick~Henry Lehmer, \emph{Interesting series involving the central binomial
  coefficient}, Amer. Math. Monthly \textbf{92} (1985), no.~7, 449--457.
  \MR{0801217 (87c:40002)}

\bibitem{EIS}
Neil J.~A. Sloane, \emph{The {O}n-{L}ine {E}ncyclopedia {O}f {I}nteger
  {S}equences}, Notices Am.\ Math.\ Soc. \textbf{50} (2003), no.~8, 912--915,
  http://www.research.att.com/\char126njas/sequences/. \MR{1992789
  (2004f:11151)}

\end{thebibliography}

\documentclass{amsart}

\usepackage{amssymb}

\begin{document}

\title[Corrigenda Am. Math. Monthly 92 (1985) 449]{Corrigenda to ``Interesting Series Involving the Central Binomial Coefficient'' [Am. Math. Monthly vol 92 (1985)]}

\author{Richard J. Mathar}
\urladdr{http://www.strw.leidenuniv.nl/~mathar}
\email{mathar@strw.leidenuniv.nl}
\address{Leiden Observatory, Leiden University, P.O. Box 9513, 2300 RA Leiden, The Netherlands}

\subjclass[2000]{Primary 40-01, 40H05; Secondary 00A20}

\date{\today}
\keywords{Errata, Central Binomial Coefficient, series}

\begin{abstract}
These are seven corrigenda to equations in the Lehmer article in
American Mathematical Monthly 92 (1985), pp 449--457, partially
reproduced in the Apelblat tables of integrals and series.
\end{abstract}

\maketitle

\section{Paper by D. H. Lehmer}
The corrigenda to equations in an article by the late
Derrick Lehmer \cite{LehmerAMM92} are:
\begin{enumerate}
\item
Multiply the right hand side of
the first equation on page 451 by two:
\[
\sum_{n=1}^\infty \frac{(-1)^{n+1}{2n\choose n}}{n4^n}= 2\log\left[\frac{\sqrt{2}+1}{2}\right].
\]
The numerical value is sequence A157699 in the Online Encyclopedia of Integer Sequences
(OEIS) \cite{EIS}.

\item
Switch the sign of the first but last term in equation (7):
\[
x\sum_{n=1}^\infty
\frac{{ 2n\choose n}}{n(n+1)}x^n = 2x \log\frac{1-\sqrt{1-4x}}{x}+\frac{\sqrt{1-4x}}{2}
-x(\log 4 -1)-\frac{1}{2}.
\]
\item
Multiply the right hand side of
an equation on page 453 by two:
\[
\sum_{m=1}^\infty \frac{(-1)^{m-1}}{m{2m\choose m}}= 2\log\left(\frac{1+\sqrt{5}}{2}\right)/\sqrt{5}.
\]
The numerical value is sequence A086466 in the OEIS \cite{EIS}.
\item
Flip a sign on one side of another equation  on page 453:
\[
\sum_{m=1}^\infty 
\frac{(-1)^{m-1}}{m^2{2m\choose m}}= 2\left\{ \log\left(\frac{\sqrt{5}+1}{2}\right)\right\}^2.
\]
The numerical value is sequence A086467 in the OEIS \cite{EIS}.
\item
Multiply the right hand side of the last line of another formula on page 453 by four:
\[
\sum_{m=1}^\infty 
\frac{1}{m^3{2m\choose m}}= -\frac{4\zeta(3)}{3}-\frac{\pi\sqrt{3}}{18}\left\{
\psi\left(\frac{1}{3}\right)-\psi\left(\frac{2}{3}\right)
\right\}.
\]
Variants in notation have appeared since then, see theorem 3.3 of a paper
by Borwein et al.\ \cite{BorweinEM10}, which 
is
equation (47) in a later work \cite{BorweinAM70},
or equation (2.67) by Davydychev and Kalmykov
\cite{DavydychevNPB699}.
The numerical value is sequence A145438 in the OEIS \cite{EIS}.
\item
Drop a factor two on the right hand side of a formula on page 455
and change sign on either side:
\[
\sum\frac{(-1)^m m^3}{{2m \choose m}}
=
\frac{2}{625}[14\sigma+5]
.
\]
The numerical value is sequence A157701 \cite{EIS}.
\item
The last equation on page 457 cannot be reproduced;
the value of the sum is $0.280851790115\ldots$ whereas $\pi^2/36=\zeta(2)/6\approx 0.27415567\ldots$
Equation (13) of page 453 yields the substitute
\[
\sum\frac{2^m(2-\sqrt{3})^m}{m^2{2m\choose m}}
=
2\left(\arcsin \tau\right)^2,\quad
\tau \equiv \frac{\sqrt{3}-1}{2} = \sqrt{2}\sin\frac{\pi}{12} =
\sin\frac{\pi}{3}-\sin\frac{\pi}{6}
.
\]
\end{enumerate}

\section{Apelblat Tables}

Equivalent corrections concern
section 4.1 of the Apelblat tables \cite{Apelblat2}:
\begin{enumerate}
\item
Insert an exponent 2 on the left hand side of formula 40:
\[
\sum_{n=1}^\infty \frac{(-1)^{n-1}n^2(n!)^2}{(2n)!} =
\frac{4}{125}\left[ 5-\sqrt{5}\ln\left(\frac{\sqrt{5}+1}{2}\right)\right]
.
\]
The numerical value is sequence A145434 \cite{EIS}.
\item
Change two numbers on the right hand side of formula 42:
\[
\sum_{n=1}^\infty \frac{(-1)^{n-1}(n!)^2}{n(2n)!} =
\frac{2}{\sqrt{5}}\ln\left(\frac{1+\sqrt{5}}{2}\right)
.
\]
\item
Multiply the right hand side of formula 47 by four and
show that the $\psi$ are trigamma functions:
\[
\sum_{n=1}^\infty \frac{(n!)^2}{n^3(2n)!} =
4\left\{
\frac{\pi\sqrt{3}}{72}\left[\psi^{(1)}\left(\frac{2}{3}\right)-\psi^{(1)}\left(\frac{1}{3}\right)\right]-\frac{1}{3}\zeta(3)
\right\}
.
\]
\item
Replace the right hand side of formula 81:
\[
\sum_{n=1}^\infty \frac{2^n(2-\sqrt{3})^n(n!)^2}{n^2(2n)!}=2\arcsin^2\left(
\frac{\sqrt{3}-1}{2}
\right)
.
\]

\end{enumerate}

\bibliographystyle{amsplain}
\bibliography{all}

\end{document}